OpenStudy (anonymous):

Precalc- Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4> 20.3° 10.2° 0.2° 30.3°

OpenStudy (anonymous):

let x be the angle between the vectors then we have cos x = (6*7 +(-1)*(-4) )/ sqrt(37*65) or x = cos-1 (46/sqrt(37*65)) or x = 43.17 appx

OpenStudy (anonymous):

@matricked are you sure thats right because these are the answer chioces 20.3° 10.2° 0.2° 30.3°

OpenStudy (anonymous):

either my calculation is wrong or the check the question...